# AutoGCN - Towards Generic Human Activity Recognition with Neural Architecture Search

### 02 Feb 2024 v1

@@ 4.2 Implementation
In the experimental setup, the student architecture is trained for a maximum of 25 epochs, while the argmax architecture is trained for 80 epochs to ensure complete convergence. The learning rate used in training is sampled from the search space and undergoes a warm-up for the first ten epochs with a decay factor of $0.5$. Subsequently, the learning rate is reduced by $0.25$ at epochs 30, 50, 60, 65, and 70. An early stopping mechanism is employed to identify underperforming student architectures during the initial training phase, terminating the training process before the 6th epoch if the accuracy is below 50%.
@@
The training is performed on a single NVIDIA-V100 with 32 GB GPU RAM on the PyTorch framework (version 1.13.0) [39] with the global seed set to $1234$. The code is publicly available under https://github.com/DeepInMotion/AutoGCN.
@@ 4.3 Comparison with other SOTA
To evaluate the performance of AutoGCN, it is compared with the baseline approach from Peng et al. [19] and other state-of-the-art (SOTA) HAR approaches. The best-performing models’ results on the NTU RGB+D 60 database are shown in Table 2 and 3, respectively. The values of $\mathcal{P}_{max}(\alpha,h)$ for the best-performing model are listed in Table 4.
@@
As shown in Table 2, our approach surpasses the NAS-GCN baseline with the Joint configuration by 0.8% for the X-View and the X-Sub dataset. In contrast to SNAS-GCN, our model demonstrates a notable increase in accuracy, with a substantial gain of 1.2% observed for both the X-Sub and X-View datasets. Moreover, it is essential to emphasize that AutoGCN achieves similar results to the baseline method NAS-GCN without the need for ensemble techniques, achieving a competitive accuracy of 95.5% on the X-View dataset. These results show the effectiveness and efficiency of AutoGCN in the context of skeleton-based human action recognition by delivering high performance without the added complexity of ensemble methods.
Table 3 demonstrates that our approach yields results comparable to the SOTA models for the NTU RGB+D 120 dataset previously reported.
As delineated in Table 4, a discernible disparity in the identified architectural components between the X-Sub and X-View datasets emerges. This disparity supports our contention from the introduction that these two datasets have inherent characteristics that cannot be effectively captured by a single searched architectural framework alone.
@@
Attention layer Ja Fa
Conv. layer type SG Bottleneck
Dropout probability 0.15 0.25
Init layer size 256 64
@@
Scaling factor 0.6 0.4
Temporal window 3 5
Graph distance 2 3
Reduction factor 1.35 1.25
@@
Depth main 3 1
Graph distance 9 11
@@
Optimizer AdamW SGD
Learning rate 0.01 0.05
Weight decay 0.0 0.001
@@ 4.4 Comparison to Random Search
Random search [45] as an optimization technique for tuning hyperparameters is widely used in NAS procedures, which makes it an essential component of our experiments [46]. To ensure the fairest comparison against our proposed method, we perform the random search following the same steps as our algorithm AutoGCN: An initial cohort of 20 student architectures are trained for 25 epochs, and subsequently, the highest-performing architecture is chosen and trained for an extended training period of 80 epochs. Finally, we compare the iterations taken by each approach to obtain the best-performing architectures, displayed in Table 5. It can be observed that the AutoGCN algorithm achieves a higher-performing architecture after just 20 iterations both for the X-View and X-Sub dataset compared to random search. To achieve a comparable point-estimates as AutoGCN, random search has to undergo another 20 iterations. The results indicate that AutoGCN can obtain an optimal architectural configuration without being subject to the stochasticity of random search.
@@ 4.5 Influence of the Controller hyperparmeters
Since the controller has fixed hyperparameters, we investigate the ideal number of rollouts for training our model, aiming to balance computational efficiency and the final model performance. The found values are shown in Table 6, in which the Top-1 accuracy percentages in relation to the rollout hyperparameter of the controller and the update stages are depicted. It also displays the average accuracies of the student models and Top-1 accuracies for the first, second, and third updates in percentage values. The table provides an overview of how the choice of rollouts impacts the accuracy of the two datasets, X-Sub and X-View. The arrow symbols describe the trend in the accuracy and average accuracy compared to the previous update: ’↑’ for an increase, ’↓’ for a decrease, ’→’ for no significant change.
@@ 4.6 Influence of the model size
We conduct experiments to investigate the optimal network size by altering the search space, as outlined in Table 1. These focus on exploring larger architectures, where we increased layers, blocks, and depth sizes in the experimental design.
@@
The results in Table 7 indicate that the X-View dataset’s accuracy remains consistent for the larger and smaller sampled architecture after 20 rollouts. The small model requires much fewer parameters but slightly more FLOPs than the larger model. The greater amount of FLOPs is due to the more complex Bottleneck convolutional layer chosen by the controller.
This outcome shows the effectiveness of the proposed search space, as defined in Table 1, and emphasizes the usage of smaller models that require significantly less training. The results advocate utilizing these smaller models for their computational efficiency while achieving the same performance as the large model, requiring fewer computational resources.
@@ 4.7 Data representation
To assess the data representation’s robustness, an examination of the effectiveness of the introduced acceleration feature’s effectiveness is conducted. Consequently, this feature is excluded and incorporated into the highest-performing model identified through our search. The impact of omitting this feature from the model becomes evident when observing the results presented in Table 8, where the Top-1 accuracy demonstrates a notable decrease. When conducting a two-sided z-test with a 5% significance level, the null hypothesis of no significant difference between those data modalities can be rejected.
@@ 5 Conclusion
In this study, we have developed a GCN NAS algorithm named AutoGCN tailored for the task of skeleton-based HAR. We formulated the intricate dependencies between hyperparameters and architectural configurations in an expressive search space encompassing a broad range of building blocks from which the controller can sample. That allows the algorithm to craft a versatile and high-performing network architecture. Accordingly, we have applied AutoGCN to identify and optimize these fundamental building blocks of the network concurrently with a reinforcement algorithm, giving every search space parameter a policy that can be updated based on the anticipated performance from the sampled student architectures. Furthermore, we have refined the search process’ exploration and exploitation behavior by incorporating a replay memory during the search process. This enables us to strike a promising balance between those, enhancing the overall search performance.
@@
Through extensive experiments, we provide a rigorous performance analysis, comparing our method against the baseline NAS procedure but also against SOTA approaches in the domain of skeleton-based HAR. Finally, AutoGCN proves its effectiveness when compared with random search.
@@
In the future, we want to investigate the influence of performance predictors on AutoGCN, which could shorten the search time of the algorithm since good-performing architectures are recognized at an earlier stage. Additionally, weight-sharing techniques could further boost the algorithm.

### 12 Mar 2024 v3

@@ 4.2 Implementation
In the experimental setup, the student architecture is trained for a maximum of 25 epochs, while the argmax architecture is trained for 80 epochs to ensure complete convergence. The learning rate used in training is sampled from the search space and undergoes a warm-up for the first ten epochs with a decay factor of $0.5$. Subsequently, the learning rate is reduced by $0.25$ at epochs 30, 50, 60, 65, and 70. An early stopping mechanism is employed to identify underperforming student architectures during the initial training phase, terminating the training process before the sixth epoch if the accuracy is below 50%.
@@
The training is performed on a single NVIDIA-V100 with 32 GB GPU RAM on the PyTorch framework (version 2.0.1) [39] with the global seed set to $1234$. The code and the experiment results are publicly available at https://github.com/DeepInMotion/AutoGCN.
@@ 4.3 Comparison with other SOTA
To evaluate the performance of AutoGCN, it is compared with the baseline approach from Peng et al. [19] and other state-of-the-art (SOTA) HAR approaches. The best-performing models’ results on the NTU RGB+D 60 database are shown in Table (2) and (3), respectively. The values of $\mathcal{P}_{max}(\alpha,h)$ for the best-performing model are listed in Table (4). Moreover, the values of the Policies $\mathcal{P}$ for all search space parameters are listed in Appendix (A) Fig. (4) and Appendix (B) Fig. (5), accompanied by an analysis of these values.
@@
As shown in Table (2), our approach surpasses the NAS-GCN baseline with the Joint configuration by 0.8% for the X-View and the X-Sub dataset. In contrast to SNAS-GCN, our model demonstrates a notable increase in accuracy, with a substantial gain of 1.2% observed for both the X-Sub and X-View datasets. Moreover, it is essential to emphasize that AutoGCN achieves similar results to the baseline method NAS-GCN without the need for ensemble techniques, achieving a competitive accuracy of 95.5% on the X-View dataset. These results show the effectiveness and efficiency of AutoGCN in the context of skeleton-based human action recognition by delivering high performance without the added complexity of ensemble methods.
Table (3) demonstrates that our approach yields results comparable to the SOTA models for the NTU RGB+D 120 dataset previously reported.
As delineated in Table (4), a discernible disparity in the identified architectural components between the X-Sub and X-View datasets emerges. This disparity supports our contention from the introduction that these two datasets have inherent characteristics that cannot be effectively captured by a single searched architectural framework alone.
@@
Attention layer Fa Ja
Conv. layer type Bottleneck SG
Dropout probability 0.25 0.15
Init layer size 64 256
@@
Scaling factor 0.4 0.6
Temporal window 5 3
Graph distance 3 2
Reduction factor 1.25 1.35
@@
Depth main 1 3
Graph distance 11 9
@@
Optimizer SGD AdamW
Learning rate 0.05 0.01
Weight decay 0.001 0.0
@@ 4.4 Comparison to Random Search
Random search [45] as an optimization technique for tuning hyperparameters is widely used in NAS procedures, which makes it an essential component of our experiments [46]. To ensure the fairest comparison against our proposed method, we perform the random search following the same steps as our algorithm AutoGCN: An initial cohort of 20 student architectures are trained for 25 epochs, and subsequently, the highest-performing architecture is chosen and trained for an extended training period of 80 epochs. Finally, we compare the iterations taken by each approach to obtain the best-performing architectures, displayed in Table (5). It can be observed that the AutoGCN algorithm achieves a higher-performing architecture after just 20 iterations both for the X-View and X-Sub datasets compared to random search. To achieve a comparable point-estimates as AutoGCN, random search has to undergo another 20 iterations. The results indicate that AutoGCN can obtain an optimal architectural configuration without being subject to the stochasticity of random search.
@@ 4.5 Influence of the Controller hyperparmeters
Since the controller has fixed hyperparameters, we investigate the ideal number of rollouts for training our model, aiming to balance computational efficiency and the final model performance. The found values are shown in Table (6), in which the Top-1 accuracy percentages in relation to the rollout hyperparameter of the controller and the update stages are depicted. It also displays the average accuracies of the student models and Top-1 accuracies for the first, second, and third updates in percentage values. The table provides an overview of how the choice of rollouts impacts the accuracy of the two datasets, X-Sub and X-View. The arrow symbols describe the trend in the accuracy and average accuracy compared to the previous update: ’↑’ for an increase, ’↓’ for a decrease, ’→’ for no significant change. Furthermore, those results are visualized in Fig. (3).
@@
Figure 3: Influence of the rollout parameter on the model performance. The star indicates the highest accuracy achieved in the rollout experiments.
@@ 4.6 Influence of the model size
We conduct experiments to investigate the optimal network size by altering the search space, as outlined in Table (1). These focus on exploring larger architectures, where we increased layers, blocks, and depth sizes in the experimental design.
@@
The results in Table (7) indicate that the X-View dataset’s accuracy remains consistent for the larger and smaller sampled architecture after 20 rollouts. The small model requires much fewer parameters but slightly more FLOPs than the larger model. The greater amount of FLOPs is due to the more complex Bottleneck convolutional layer chosen by the controller.
This outcome shows the effectiveness of the proposed search space, as defined in Table (1), and emphasizes the usage of smaller models that require significantly less training. The results advocate utilizing these smaller models for their computational efficiency while achieving the same performance as the large model, requiring fewer computational resources.
@@ 4.7 Data representation
The newly introduced acceleration feature is analyzed to gauge the strength of the data representation. Consequently, this feature is excluded and incorporated into the highest-performing model identified through our search. The impact of omitting this feature from the model becomes evident when observing the results presented in Table (8), where the Top-1 accuracy demonstrates a notable decrease. When conducting a two-sided z-test with a 5% significance level, the null hypothesis of no significant difference between those data modalities can be rejected.
@@
5 Limitations & Future Work
The proposed AutoGCN algorithm uses a reinforcement learning approach, requiring every student network to be trained from scratch at each iteration. While this strategy enables dynamic adaptation and refinement of the individual architecture and hyperparameter components, it also leads to an increased computational burden in computing the reward for the controller guiding the search and updating the policies $\mathcal{P}$.
In order to avoid having to train every student architecture from scratch, the search procedure could adopt so-called one-shot techniques like DARTs, in which an over-parameterized supernetwork is trained to contain every potential architecture build and enable weight sharing among students [12, 15]. To implement this approach, reusable building blocks, and operation types that can be recycled among different student compositions effectively would have to be defined. While such an adaptation could then lead to a speed-up of the search procedure, it could also lower the versatility of the search space.
In our work, speed-up techniques for the search process are limited to leveraging the knowledge reservoir $\mathcal{N}$ and implementing an early-stopping mechanism. Possible enhancements to accelerate the search could involve the inclusion of performance predictors [12]. Such predictors could assess the potential of a given student architecture early, allowing for a more efficient allocation of computational resources by prioritizing the exploration of promising candidates and neglecting unpromising ones. With such techniques, the convergence towards optimal solutions could be accelerated, resulting in a significantly reduced overall search time.
@@ 6 Conclusion
In this study, we have developed and presented a GCN NAS algorithm named AutoGCN, tailored for the task of skeleton-based HAR. The intricate dependencies between hyperparameters and architectural configurations are formulated in an expressive search space encompassing a broad range of building blocks from which the controller can sample, which allows the algorithm to find a versatile and high-performing network architecture. AutoGCN is applied to identify and optimize these fundamental building blocks of the network concurrently with a reinforcement algorithm, giving every search space parameter a policy that can be updated based on the anticipated performance from the sampled student architectures. The search process’s exploration and exploitation behavior is refined by incorporating a replay memory during the search process, enabling the method to strike a promising balance between these and enhancing the overall search performance.
@@
Through extensive experiments, we provide a rigorous performance analysis, comparing our method against the baseline NAS procedure and SOTA approaches in the domain of skeleton-based HAR. Finally, AutoGCN demonstrates effectiveness also compared to random search.
@@
Future work will investigate the influence of performance predictors on AutoGCN to decrease the search time of the algorithm, as good-performing architectures could be recognized earlier. Additionally, weight-sharing techniques could further make the search more efficient.
Appendix A Policy values on the X-Sub dataset
Fig. (4) shows the policy values for the model’s hyperparameters and architecture search space with 20 rollouts between controller updates on the X-Sub dataset. These values are categorized within the four search areas as shown in Table (1). The highest Top-1 accuracy in this experiment is attained with the first controller update, with the final $\mathcal{P}_{max}(\alpha,h)$ values outlined in Table (4).
Within the Common group, most policy values maintain consistency after the second update cycle. However, within the Conv. layer choices, uncertainty arises, particularly evident in the fluctuation of probabilities between the Basic and Shuffle layers after the first controller update. A similar trend of uncertainty is visible in the Dropout probability and Init layer size. Regarding the Input stream group, discernible trends emerge after both the first and second update cycles. The search space parameters with lower probability values tend to retain their trend after the first update. Nonetheless, the controller is uncertain regarding Blocks in, Stride in, and the Reduction factor, with a tight distribution across these search space parameters for the highest values. In the Main stream group, most search areas demonstrate clear trends, barring the values associated with Depth main. These probability values for the depths of one or three layers remain closely clustered following the first and second controller updates. Within the Optimizer group, variability among the policy values is notable, particularly for the Learning rate and Batch size parameters. Given the intertwined nature of these hyperparameters and their impact on the incompletely trained student architecture, fluctuations are expected.
It is essential to highlight that multiple local optima are possible within this search space configuration. These optima depend upon the statistical variance from the random sampling process of the student architectures. Furthermore, the average student’s Top-1 accuracy is significantly lower with the second update cycle in this experiment compared to the other rollout experiments from Table (6).
Figure 4: Policy values for the best-performing model on the X-Sub dataset. One update cycle contains 20 student architectures.
Appendix B Policy values on the X-View dataset
Fig. (5) shows the policy values for the model’s hyperparameter and architecture search space with 30 rollouts between controller updates on the X-View dataset. The values are grouped as defined in Table (1). The best Top-1 accuracy in this experiment is achieved with the second controller update, where the $\mathcal{P}_{max}(\alpha,h)$ values are presented in Table (4).
For the Common group, the trend from the first to the second controller update is relatively consistent in its upward direction for the highest search space value. In contrast, the final value for the Init layer size is experiencing a significant increase in its probability value. The policy values with lower probabilities fluctuate more between the first and second controller updates. In the Input stream group, most policy values indicate a clear trend for the one single search space value. In contrast, the Graph distance and the Reduction factor show variability between the first and the second update. Regarding the Main stream group, most search areas manifest clear trends, except for the values associated with Depth main. As in the X-Sub experiment, the probability values for depths of one or three layers remain closely clustered following the first and second controller updates. Additionally, the values for the Residual layer converse after the second update. In the Optimizer group, the policy values indicate less fluctuation than for the X-Sub dataset in Fig. (4). A clear trend among these different groups is visible from the first to the second controller update.
As with the X-Sub experiments, the search space values contain a statistical variance due to the random sampling process of the student architectures.
Figure 5: Policy values for the best-performing model on the X-View dataset. One update cycle contains 30 student architectures.